OpenStudy (anonymous):

If x=cost, y=cos2t then the equation of the curve y=f(x) is : a)y=2x^2-1 b)y=x^2+1 c)y=1-2x^2 d)y=x^2/2 -1

ganeshie8 (ganeshie8):

hint : cos2t = 2cos^2t - 1

OpenStudy (anonymous):

@ganeshie8 i got y+2costsint=2x(x-cost)

ganeshie8 (ganeshie8):

\(y = \cos 2t = 2\cos^2t-1 = 2x^2-1\)

OpenStudy (anonymous):

oh it says the equation of the curve i didn't c it well i thought it want the tangent equation l so i found the derivative by chain rule and it got complicated LOL

OpenStudy (anonymous):

thanx alot :D